# Formal Analysis of the Contract Automata Runtime Environment with Uppaal: Modelling, Verification and Testing

### v1

@@ 7.1. Testing the Orchestrator
Different variants of the model have been developed, each one equipped with specific test code to test one of the components. One version of the model is equipped with test code for testing the orchestrator (i.e., RunnableOrchestration). In this case, during concrete testing, the orchestrator will be the actual system, whilst the tester will mimic the services. Therefore, the test code will be generated from the services model. In this version, the automaton template RunnableOrchestratedContract contains test code in its transitions, whilst the automaton template RunnableOrchestration does not contain test code. However, useful comments are inserted as text code in some transitions of RunnableOrchestration to guide the generation of concrete tests from abstract tests. This mainly involves printing the actual value during simulation of variables offerer, requester and i used by the RunnableOrchestration as indices to identify the services the orchestrator is interacting with. In the generated tests, no timeout is experienced, therefore the SocketTimeoutAutomaton is a dummy automaton receiving signals from RunnableOrchestratedContract and doing nothing.
@@ 7.3. An example of test generation
We now discuss a query used for testing an orchestration whose set-up is dictatorial choice with centralised action. This means that the orchestrator will decide each choice and the orchestrator will act as a broker in a match between the two services. All versions of the model contains queries for testing the various configurations of the orchestrator (e.g., majoritarian choice, distributed action). We refer to [6] for more details about these other queries. In the concrete tests, the orchestrator will interact with two mocked services. Furthermore, the orchestration contract automaton, a parameter of the RunnableOrchestration object (see Figure 2), is dummy and is instrumental to cover the desired portions of code of CARE. In particular, the orchestration contract automato that will be used in the concrete tests accepts the trace (!euro,-)(?coffee,!coffee)[(!euro,-)]. The corresponding query used to generate the abstract test is:
@@
alice.steps[2]==CENTRALISED_OFFER &&
alice.steps[3]==CENTRALISED_MATCH &&
bob.steps[4]==CENTRALISED_OFFER &&
@@
alice.steps[7]==CENTRALISED_OFFER &&
@@
The query above simulates a run accepted by the orchestration contract automaton. We now analyse the query. As discussed in Section 5, the orchestration always starts by checking if the configurations of the services and the orchestrator are matching. Therefore, both services are instructed to check their configuration (alice.steps[0]==ORC_CHECK and bob.steps[1]==ORC_CHECK). In this test, the configuration is of centralised action and dictatorial choice. After the configuration checks, the orchestration starts. In the query, Alice is instructed to execute the offer action !euro with alice.steps[2]==CENTRALISED_OFFER. This corresponds to execution of the offer label (!euro,-) in the orchestration (Alice is the first service). After that, the match label (?coffee,!coffee) of the orchestration is executed. The match starts with Alice performing the request ?coffee, encoded in the query as alice.steps[3]==CENTRALISED_MATCH. After Alice performs the request action, it is Bob’s turn to match the request of Alice with the corresponding offer action!coffee. In Figure 9, it is possible to observe that the transition of the RunnableOrchestratedContract Uppaal template covering the execution of an offer action in both the cases of offer label or match label (respectively, d1==NOPAYLOAD or d1==REQUEST, see Section 5) has in its effect the instruction steps[step]=CENTRALISED_OFFER. Therefore, Bob is instructed to execute the offer action !coffee in the query with bob.steps[4]==CENTRALISED_OFFER.
@@
We now discuss the abstract test code generated by the query and how it is concretised into a JUnit test for the real system. Uppaal also allows the insertion of abstract test code in the system declaration of a model, which will be printed out in the abstract test either at the start (i.e., pre-fix) or upon termination of the test (i.e., post-fix). In all tests, pre-fix test code is used to declare some variables that will be utilised during the test. The first transitions executed during the test are those related to the configuration check. Both Alice and Bob generate the same portion of abstract test code, reported below.
@@
[mathescape,fontsize=,linenos]java @Test public void test() throws IOException, ClassNotFoundException, InterruptedException alice = new Thread(()-¿ check(8080); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8080)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_alice(oin,oout);
@@
bob = new Thread(()-¿ check(8081); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8081)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_bob(oin,oout); catch (IOException — ClassNotFoundException e) throw new RuntimeException(e); ); alice.start(); bob.start(); RunnableOrchestration ror = new DictatorialChoiceRunnableOrchestration(new Agreement(), adc.importMSCA(dir+”Orchestration.data”), Arrays.asList(null,null), Arrays.asList(8080,8081), new CentralisedOrchestratorAction());
@@
After the orchestration check has been performed, the orchestration starts, and the offer label (!euro,-) of the orchestration will be executed. The following instructions are appended to the abstract test case being generated.
@@
The transition outgoing state CentralisedOffer of the RunnableOrchestration automaton (see Figure 6) has as test code //centralised offerer \$(ror.offerer). This will append the comment at line 1 in the abstract test, indicating that the next instructions of the abstract test will cover the first offer action performed by Alice (service at index 00). After performing the check, both Alice and Bob are again in state Ready (Figure 9). The first readObject operation (line 3) is again appended by the outgoing transition from state Ready of Alice. The next transition executed by Alice, with guard d1==ACTION, will append the instruction at line 5. The assertion will check whether the message received by the orchestrator will be the expected action, which will be concretised in the concrete test with the value "euro" (i.e., the orchestrator is invoking the method euro of Alice).
@@
This concludes the test code for the first offer label (!euro,-) of the orchestration. Next, the match label (?coffee,!coffee) is executed. The instructions appended to the abstract test case during the execution of the match label are reported below.
@@
This concludes the match (?coffee,!coffee). As stated earlier, the orchestrator will now perform a choice on whether to continue and execute the offer label (!euro,-) or terminate. For brevity, we only report the abstract test code dumped by a service when executing a dictatorial choice, reported below.
@@
The instruction at line 1 is again appended when executing the dequeue effect outgoing state Ready (Figure 9). After that, the transition with guard d1 == ORC_CHOICE is executed, which will append the test code at line 3. The next executed transition, with effect steps[step]=DICTATORIAL_CHOICE, has no test code generated but it is used to fulfill the query. No other test code is generated for the dictatorial choice, as the orchestrator decides on its own the next move. This test code is generated by both Alice and Bob, performing in turn the choice. In this example, after the first dictatorial choice, the orchestrator will execute again the offer label (!euro,-), and again a dictatorial choice. We have already covered the test code that will be generated. After that, the orchestrator decides to terminate. The test code generated by each service is:
@@ 7.4. Coverage
The generated abstract tests cover all transitions of the model and all interactions between the orchestrator and the services. The code coverage indicates that the concrete tests derived from the model cover a significant portion of the source code. This suggests that the model is not excessively abstract compared to the actual implementation.
@@ 8. Conclusion and Future Work
We have presented the formal modeling and verification of the Contract Automata Runtime Environment (CARE), an open-source platform. Both statistical and exhaustive model checking techniques have played a crucial role in formally verifying numerous desired properties of the modeled system, such as the absence of deadlocks, while also enhancing the accuracy of the formal model. Statistical model checking has been employed to fine-tune parameter settings within the formal model, such as the buffer size. The adequacy of the abstract model has been described, and the transitions of the formal model have been linked to the corresponding lines of source code. The tests generated from the formal model have been employed to test the source code.
At present, the different artifacts such as the model, source code, tests, and tracing information are manually kept aligned. This process demands substantial effort whenever a new version of CARE is introduced, as each artifact needs to be updated accordingly. Future work involves studying techniques for automatic alignment of these artifacts. Furthermore, it would be interesting to exploit the facilities provided by the recently introduced tool Uppex [43] to factorize the different variants of the model discussed in the paper under a single, configurable Uppex model, and to automate the selection of a configuration.

### v2

@@ 7.1. Testing the Orchestrator
Different variants of the model have been developed, each one equipped with specific test code to test one of the components. One version of the model is equipped with test code for testing the orchestrator (i.e., RunnableOrchestration). In this case, during concrete testing, the orchestrator will be the actual system, whilst the tester will mimic the services. Therefore, the test code will be generated from the services model. In this version, the automaton template RunnableOrchestratedContract contains test code in its transitions, whilst the automaton template RunnableOrchestration does not contain test code. However, useful comments are inserted as test code in some transitions of RunnableOrchestration to guide the generation of concrete tests from abstract tests. This mainly involves printing the actual value during simulation of variables offerer, requester and i used by the RunnableOrchestration as indices to identify the services the orchestrator is interacting with. In the generated tests, no timeout is experienced, therefore the SocketTimeoutAutomaton is a dummy automaton receiving signals from RunnableOrchestratedContract and doing nothing. Indeed, a timeout in the concrete system could be experienced due to a worsening of the physical network, which is outside the scope of the model-based testing performed in this paper to ascertain the adequacy.
@@ 7.3. An example of test generation
We now discuss an example of testing an orchestration whose set-up is dictatorial choice with centralised action. This means that the orchestrator will decide each choice and the orchestrator will act as a broker in a match between the two services.
Query
All versions of the model contains queries for testing the various configurations of the orchestrator (e.g., majoritarian choice, distributed action). We refer to [10] for more details about these other queries. In the concrete tests, the orchestrator will interact with two mocked services. Furthermore, the orchestration contract automaton, a parameter of the RunnableOrchestration object (see Figure 3), is dummy and is instrumental to cover the desired portions of code of CARE. In particular, the orchestration contract automaton that will be used in the concrete tests accepts the trace (!euro,-)(?coffee,!coffee)[(!euro,-)]. The corresponding query used to generate the abstract test is:
@@
alice.steps[2]==CENTRALISED_OFFER && //alice offer action !euro
alice.steps[3]==CENTRALISED_MATCH && //alice request action ?coffee
bob.steps[4]==CENTRALISED_OFFER && //bob offer action !coffee
@@
alice.steps[7]==CENTRALISED_OFFER && //alice offer action !euro
@@
The query above is endowed with comments to pinpoint each step with the corresponding action, if any. An evidence to the query above simulates a run accepted by the orchestration contract automaton. We now analyse the query. As discussed in Section 5, the orchestration always starts by checking if the configurations of the services and the orchestrator are matching. Therefore, both services are instructed to check their configuration (alice.steps[0]==ORC_CHECK and bob.steps[1]==ORC_CHECK). In this test, the configuration is of centralised action and dictatorial choice. After the configuration checks, the orchestration starts. In the query, Alice is instructed to execute the offer action !euro with alice.steps[2]==CENTRALISED_OFFER. This corresponds to execution of the offer label (!euro,-) in the orchestration (Alice is the first service). After that, the match label (?coffee,!coffee) of the orchestration is executed. The match starts with Alice performing the request ?coffee, encoded in the query as alice.steps[3]==CENTRALISED_MATCH. After Alice performs the request action, it is Bob’s turn to match the request of Alice with the corresponding offer action!coffee. In Figure 12, it is possible to observe that the transition of the RunnableOrchestratedContract Uppaal template covering the execution of an offer action in both the cases of offer label or match label (respectively, d1==NOPAYLOAD or d1==REQUEST, see Section 5) has in its effect the instruction steps[step]=CENTRALISED_OFFER. Therefore, Bob is instructed to execute the offer action !coffee in the query with bob.steps[4]==CENTRALISED_OFFER.
@@
7.3.1. Test generation
We now discuss the abstract test code generated by the query and how it is concretised into a JUnit test for the real system. Uppaal also allows the insertion of abstract test code in the system declaration of a model, which will be printed out in the abstract test either at the start (i.e., pre-fix) or upon termination of the test (i.e., post-fix). In all tests, pre-fix test code is used to declare some variables that will be utilised during the test.
Configuration check. The first transitions executed during the test are those related to the configuration check. Both Alice and Bob generate the same portion of abstract test code, reported below.
@@ 7.3.1. Test generation
{listing}
The JUnit method testing the centralised dictatorial orchestrator
@@
[mathescape,fontsize=,linenos]java @Test public void test() throws IOException, ClassNotFoundException, InterruptedException Thread alice = new Thread(()-¿ //mocked service check(8080); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8080)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_alice(oin,oout);
@@
Thread bob = new Thread(()-¿ //mocked service check(8081); try (ServerSocketChannel server = ServerSocketChannel.open()) server.bind(new InetSocketAddress(8081)); Socket socket = server.accept().socket(); ObjectInputStream oin = new ObjectInputStream(socket.getInputStream()); ObjectOutputStream oout = new ObjectOutputStream(socket.getOutputStream()); oout.flush(); uppaal_bob(oin,oout); catch (IOException — ClassNotFoundException e) throw new RuntimeException(e); ); alice.start(); bob.start(); RunnableOrchestration ror = new DictatorialChoiceRunnableOrchestration(new Agreement(), adc.importMSCA(dir+”Orchestration.data”), Arrays.asList(null,null), Arrays.asList(8080,8081), new CentralisedOrchestratorAction());
@@
{listing}
The JUnit method testing the configuration of the service
@@
Offer label. After the orchestration check has been performed, the orchestration starts, and the offer label (!euro,-) of the orchestration will be executed. The following instructions are appended to the abstract test case being generated.
@@
The transition outgoing state CentralisedOffer of the RunnableOrchestration automaton (see Figure 9) has as test code //centralised offerer \$(ror.offerer). This will append the comment at line 1 in the abstract test, indicating that the next instructions of the abstract test will cover the first offer action performed by Alice (service at index 00). This is helpful for generating the concrete test from the abstract test. After performing the check, both Alice and Bob are again in state Ready (Figure 12). The first readObject operation (line 3) is again appended by the outgoing transition from state Ready of Alice. The next transition executed by Alice, with guard d1==ACTION, will append the instruction at line 5. The assertion will check whether the message received by the orchestrator will be the expected action, which will be concretised in the concrete test with the value "euro" (i.e., the orchestrator is invoking the method euro of Alice).
@@
This concludes the test code for the first offer label (!euro,-) of the orchestration.
Match label. Next, the match label (?coffee,!coffee) is executed. The instructions appended to the abstract test case during the execution of the match label are reported below.
@@
This concludes the match (?coffee,!coffee).
Choice. As stated earlier, the orchestrator will now perform a choice on whether to continue and execute the offer label (!euro,-) or terminate. For brevity, we only report the abstract test code dumped by a service when executing a dictatorial choice, reported below.
@@
The instruction at line 1 is again appended when executing the dequeue effect outgoing state Ready (Figure 12). After that, the transition with guard d1 == ORC_CHOICE is executed, which will append the test code at line 3. The next executed transition, with effect steps[step]=DICTATORIAL_CHOICE, has no test code generated but it is used to fulfill the query. No other test code is generated for the dictatorial choice, as the orchestrator decides on its own the next move. This test code is generated by both Alice and Bob, performing in turn the choice. In this example, after the first dictatorial choice, the orchestrator will execute again the offer label (!euro,-), and again a dictatorial choice. We have already covered the test code that will be generated.
Termination. After that, the orchestrator decides to terminate. The test code generated by each service is:
@@
{listing}
The JUnit method testing the Alice service
@@ 7.4. Coverage
The code coverage indicates that the concrete tests derived from the model cover a significant portion of the source code. This suggests that the model is not excessively abstract compared to the actual implementation. Furthermore, from the queries we know that the generated abstract tests cover all transitions of the model and all interactions between the orchestrator and the services.
@@ 8. Conclusion and Future Work
We have presented the formal modelling and verification of the Contract Automata Runtime Environment (CARE), an open-source platform. Both statistical and exhaustive model checking techniques have played a crucial role in formally verifying numerous desired properties of the modeled system, such as the absence of deadlocks, while also enhancing the accuracy of the formal model. Statistical model checking has also been employed to fine-tune parameter settings within the formal model, such as the buffer size. The adequacy of the abstract model has been described, and the transitions of the formal model have been linked to the corresponding lines of source code. The tests generated from the formal model have been employed to test the source code.
At present, the different artifacts such as the model, source code, tests, and tracing information are manually kept aligned. This process demands substantial effort whenever a new version of CARE is introduced, as each artifact needs to be updated accordingly. Future work involves studying techniques for automatic alignment of these artifacts. Furthermore, it would be interesting to exploit the facilities provided by the recently introduced tool Uppex [59] to factorize the different variants of the model discussed in the paper under a single, configurable Uppex model, to automate the selection of a configuration and more generally to facilitate the collaboration between the software developer and the Uppaal modeller, whenever they are different persons.